# Beyond Autoregression: Discrete Diffusion for Complex Reasoning and Planning

### v1

@@ Abstract
Autoregressive language models, despite their impressive capabilities, struggle with complex reasoning and long-term planning tasks. We introduce discrete diffusion models as a novel solution to these challenges. Through the lens of subgoal imbalance, we demonstrate how diffusion models effectively learn difficult subgoals that elude autoregressive approaches. We propose Multi-granularity Diffusion Modeling (MDM), which prioritizes subgoals based on difficulty during learning. On complex tasks like Countdown, Sudoku, and Boolean Satisfiability Problems, MDM significantly outperforms autoregressive models without using search techniques. For instance, MDM achieves 91.5% and 100% accuracy on Countdown and Sudoku, respectively, compared to 45.8% and 20.7% for autoregressive models. Our work highlights the potential of diffusion-based approaches in advancing AI capabilities for sophisticated language understanding and problem-solving tasks.
@@ 1 Introduction
Building upon these insights, we propose a natural extension to current discrete diffusion models, which we term multi-granularity diffusion modeling (MDM; §3.3). This approach prioritizes different subgoals based on their difficulty during the learning process, leading to more effective learning outcomes and faster convergence.
In our experimental evaluation (§4), we focus on substantially more complex problem-solving tasks, such as Countdown (Gandhi et al., 2024) and Sudoku (Garns, 1979). These tasks demand extensive planning over a large number of combinations and pose challenges even for commercial Large Language Models (e.g., GPT-4 Achiam et al. 2023). Notably, without employing any search techniques, MDM achieves 91.5% and 100% accuracy on Countdown and Sudoku respectively, while its autoregressive counterpart only solves 45.8% and 20.7% of the problems. Additionally, we conduct experiments on the Boolean Satisfiability Problem (SAT), an NP-complete problem (Cook, 1971) that represents a wide range of constraint satisfaction problems. Our model exhibits superior performance in solving SAT problems with higher accuracy compared to the autoregressive alternative, particularly when dealing with an increased number of variables and constraints. Through this systematic exploration, we aim to demonstrate the potential advantages of diffusion-based approaches in addressing sophisticated language understanding and generation challenges. All associated code is available at https://github.com/HKUNLP/diffusion-vs-ar.
@@ 3 Subgoal Imbalance and Multi-granularity Diffusion Models
In this section, we employ a motivation example (§3.1) to elucidate the challenges faced by autoregressive models in specific scenarios. Through this analysis, we introduce the concept of subgoal imbalance—wherein some subgoals are inherently more difficult than others—which offers insights into these difficulties. We then extend our discussion in §3.2 to examine how diffusion models can more effectively address and learn these hard subgoals, effectively overcoming the limitations observed in autoregressive approaches. We finally propose Multi-granularity Diffusion Modeling (MDM; §3.3) as a natural extension of discrete diffusion models to better address these challenges and improve performance on complex tasks requiring planning and reasoning.
@@ 3.1 Subgoal Imbalance in Autoregressive and Diffusion Modeling
(Subgoal imbalance due to the unknown data distribution $q({\mathbf{x}})$) The autoregressive modeling decomposes the model distribution $p_{\bm{\theta}}({\mathbf{x}})$ into multiple subgoals: $p_{\bm{\theta}}({\bm{x}}_{1}),\dots,p_{\bm{\theta}}({\bm{x}}_{N}\mid{\bm{x}}_{1:N-1})$. Given that the true data distribution $q({\mathbf{x}})$ is unknown, the generation of individual tokens may not inherently follow an autoregressive pattern (i.e., ${\bm{x}}_{n}\not\sim q({\bm{x}}_{n}\mid{\bm{x}}_{1:n-1})$). Consequently, the difficulty of learning various subgoals can differ significantly. Given only the left context, some subgoals may require substantially more data to learn or may even be infeasible to learn.
@@ 3.2 Effective Hard Subgoal Learning in Diffusion Modeling
Further analysis of the unweighted loss $u({\mathbf{x}}_{0},{\mathbf{x}}_{t},n;{\bm{\theta}})$ in the diffusion model, based on 1,000 samples of ${\mathbf{x}}_{t}\sim q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$, reveals a clear trend: as the number of timesteps increases, resulting in more noise in ${\mathbf{x}}_{t}$, objectives in smaller timesteps (i.e., recovery from less noisy data) become significantly easier to learn. From a multi-view learning perspective (Xu et al., 2013), each ${\mathbf{x}}_{t}$ can be interpreted as a distinct view of ${\mathbf{x}}_{0}$. In the diffusion process, by exploring the consistency and complementary properties of different views offered by a diverse range of interrelated objectives $u({\mathbf{x}}_{0},{\mathbf{x}}_{t},n;{\bm{\theta}})$, our findings suggest that objectives challenging to learn in AR models become more effective, promising, and exhibit better generalization in diffusion models.
@@ 3.3 Multi-granularity Diffusion Modeling
These observations provide valuable insights, suggesting that due to subgoal imbalance, we should prioritize different subgoals based on their difficulty during the learning process to achieve more effective learning outcomes and faster convergence. Building on this, we propose the multi-granularity diffusion model as a natural extension of the discrete diffusion model.
@@
For a sequence of length $N$, the probability of sampling the same ${\mathbf{x}}_{t}$ in AR is 1. However, in diffusion, this probability reduces to $1/C_{N-1}^{t(N-1)/T}$ due to the randomness in sampling ${\mathbf{x}}_{t}$, potentially reducing the training efficiency of diffusion models. We note that Equation (7) employs a sequence-level reweighting term $w(t)$ to indicate the importance of ${\mathbf{x}}_{t}$. However, individual tokens within the sequence, given their imbalanced difficulties, are not properly reweighted. To address this, we propose multi-granularity diffusion modeling (MDM), which introduces an additional token-level reweighting mechanism to enhance training efficiency:
$L_{\mathrm{\textsc{MDM}}}=\sum_{n=1}^{N}\sum_{t=1}^{T}w(t)v({\mathbf{x}}_{t,n})u({\mathbf{x}}_{0},{\mathbf{x}}_{t},n;{\bm{\theta}}),$ (8)
@@ 4.1 Countdown
MDM 76.0 x1
As shown in Table 4.1, the performance of the GPT-4 with IO, CoT, and CoT-SC prompting methods from  Yao et al. (2024) is unsatisfactory for the given task, with only accuracy below 10%. The introduction of ToT, which incorporates a search algorithm designed by human experts into the decoding process, significantly enhances the performance of GPT-4. This integration allows the AR model to backtrack as needed, resulting in notable improvements. However, this paradigm requires the assessment of intermediate steps using LLM, resulting in considerable computational costs due to the need for multiple LLM calls. We list the token cost in Table 4.1 with more details in Appendix D.1. ToT consumes 186 times more tokens than MDM, showcasing the ‘internal’ search capability by promoting global consistency in diffusion modeling. In summary, our model, despite having a parameter size of only 85M, significantly outperforms both the AR task-specific model of the same size (GPT-2 Scratch) in terms of performance and the larger general pre-trained model (GPT-4) in computation cost, indicating it is challenging for model scaling and decoding strategies to substitute the advantages of modeling paradigm.
@@ 4.3 Boolean Satisfiability Problem
As shown in the right figure of Figure 4, MDM performs well in solving scenarios with five variables, while the AR model falls slightly short. As the number of variables increases, both our model and the AR model experience a certain degree of decrease in accuracy. However, the performance gap between the two models widens as the difficulty of the task increases. This indicates that our diffusion model exhibits a more pronounced advantage in handling more challenging tasks than the AR counterpart.
@@ 4.4 Analysis
Figure 5: Evaluation accuracy throughout the training process for AR and MDM with different reweighting strategies.
@@
We assess the trade-off between accuracy and decoding speed by comparing the performance of the AR model (GPT-2 Scratch 85M) and MDM (85M). The speed metric is determined by the number of samples processed using a batch size of 1 on the NVIDIA GeForce V100 GPU. As shown in Figure 6(a), MDM can flexibly control the trade-off between accuracy and decoding speed by varying the diffusion timesteps. Notably, by employing just one diffusion step, MDM demonstrates a remarkable 10x improvement in speed compared to AR, while maintaining superior accuracy with 75% and 12.7% compared to 45.8% and 5.1% of AR on Countdown 4 and 5, respectively. We observed that the slope of countdown 4 is smaller compared to countdown 5 in the trade-off. This suggests that for tasks with lower complexity, diffusion demonstrates a more noticeable speed advantage by setting a smaller diffusion step. In addition, it also indicates that sacrificing some efficiency for performance improvements becomes particularly evident when dealing with more intricate tasks.
@@
To gain a deeper understanding of error patterns in AR and MDM, we conducted an error analysis on Countdown 4. For instance, given the input “7,38,3,1” and the target number 14, a correct solution would be “97-38=59,59-17=42,42/3=14”. We divide the solution into four parts: equation 1, equation 2, equation 3, and answer checking. First, from a calculation perspective, we assess the error ratio in each equation by comparing the left-hand side and the right-hand side, regardless of whether the correct number was chosen. As shown in Figure 6(b), the bar plot demonstrates that the majority of calculation errors for AR are concentrated in Equation 3. For example, given input “16,4,40,51” and target 87, the prediction of AR is “51-40=11,16*4=64,11+64=87” while the correct solution is “16/4=4,40+51=91,91-4=87”. We can observe that in the last equation, the model realizes that it needs to output 87 as the final result. Therefore, even though the model most likely knows that 11 + 64 actually equals 75 given the low calculation error in the former equations, it reluctantly outputs 87 due to it being the last equation. This significantly increases the number of calculation errors in the third equation. We call this phenomenon ‘The Regretful Compromise’. The reason for this is that the AR model made incorrect choices of numbers or operations in previous equations. This is demonstrated by examining the step at which the models fail the task, as depicted in Figure 6(b). It is evident that there is a notable frequency of planning errors in the first equation for the AR model, where the number of errors is significantly higher compared to our model. This highlights the limitations of the left-to-right decoding approach in AR, which adversely affects its planning ability.
@@ 5.2 Non-autoregressive Modeling
The non-autoregressive (NAR) generation method, which produces all target tokens simultaneously given the source context, is first proposed by  (Gu et al., 2017) in the text field for machine translation, primarily due to the efficiency consideration. While a series of advancements have been made afterward (Lee et al., 2018; Gu et al., 2019; Ghazvininejad et al., 2019; Qian et al., 2021; Huang et al., 2022, inter alia), traditional NAR models still tend to underperform AR models in terms of generation quality (Xiao et al., 2023). Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), a powerful class of generative models known for their impressive image-generation capabilities (Dhariwal & Nichol, 2021), have recently been applied to the field of text generation (Hoogeboom et al., 2021; Austin et al., 2021; Li et al., 2022; Campbell et al., 2022; Dieleman et al., 2022; Chen et al., 2023; Ye et al., 2023b; Lovelace et al., 2024), reinforcement learning (Janner et al., 2022; Chi et al., 2023) and protein design (Xu et al., 2022; Hoogeboom et al., 2022b; Corso et al., 2023). In essence, diffusion models perform a multi-step denoising process to progressively convert a random noise into a data sample, and the denoising procedure can be seen as parameterizing the gradients of the data distribution (Song & Ermon, 2019), connecting them to score matching (Hyvärinen & Dayan, 2005) and energy-based models (LeCun et al., 2006). For text, the diffusion model can be seen as an extension of the traditional iterative NAR models (Gong et al., 2022) and has been shown to approach or outperform AR models on text perplexity (Han et al., 2023; Lou et al., 2023; Gulrajani & Hashimoto, 2024), diversity (Gong et al., 2022; 2023; Zhang et al., 2023) as well as various seq-to-seq tasks (Wu et al., 2023b; Zheng et al., 2023; Ye et al., 2024). In contrast, we compare diffusion with AR from a perspective of subgoal imbalance and demonstrates the effectiveness of diffusion in tasks requiring complex reasoning and planning.
@@ 6 Conclusion
This paper presents an extensive analysis of the limitations of auto-regressive (AR) language models when applied to planning tasks that involve deliberate planning, both in controlled settings and real-world contexts. Based on an advanced understanding, we propose an improved diffusion model, named MDM, that performs significantly better than AR and previous diffusion models on various sophisticated planning tasks. Our findings underscore the necessity to reevaluate the sequence modeling paradigm for modern large language models, especially in tackling challenging problem-solving tasks.

### v2

@@ Abstract
Autoregressive language models, despite their impressive capabilities, struggle with complex reasoning and long-term planning tasks. We introduce discrete diffusion models as a novel solution to these challenges. Through the lens of subgoal imbalance, we demonstrate how diffusion models effectively learn difficult subgoals that elude autoregressive approaches. We propose Multi-Granularity Mask Diffusion Modeling (MGDM), which prioritizes subgoals based on difficulty during learning. On complex tasks like Countdown, Sudoku, and Boolean Satisfiability Problems, MGDM significantly outperforms autoregressive models without using search techniques. For instance, MGDM achieves 91.5% and 100% accuracy on Countdown and Sudoku, respectively, compared to 45.8% and 20.7% for autoregressive models. Our work highlights the potential of diffusion-based approaches in advancing AI capabilities for sophisticated language understanding and problem-solving tasks. All associated codes are available at https://github.com/HKUNLP/diffusion-vs-ar.
@@ 1 Introduction
Building upon these insights, we propose a natural extension to current discrete diffusion models, which we term multi-granularity diffusion modeling (MGDM; §3.3). This approach prioritizes different subgoals based on their difficulty during the learning process, leading to more effective learning outcomes and faster convergence.
In our experimental evaluation (§4), we focus on substantially more complex problem-solving tasks, such as Countdown (Gandhi et al., 2024) and Sudoku (Garns, 1979). These tasks demand extensive planning over a large number of combinations and pose challenges even for commercial Large Language Models (e.g., GPT-4 Achiam et al. 2023). Notably, without employing any search techniques, MGDM achieves 91.5% and 100% accuracy on Countdown and Sudoku respectively, while its autoregressive counterpart only solves 45.8% and 20.7% of the problems. Additionally, we conduct experiments on the Boolean Satisfiability Problem (SAT), an NP-complete problem (Cook, 1971) that represents a wide range of constraint satisfaction problems. Our model exhibits superior performance in solving SAT problems with higher accuracy compared to the autoregressive alternative, particularly when dealing with an increased number of variables and constraints. Through this systematic exploration, we aim to demonstrate the potential advantages of diffusion-based approaches in addressing sophisticated language understanding and generation challenges.
@@ 3 Subgoal Imbalance and Multi-granularity Diffusion Models
In this section, we employ a motivation example (§3.1) to elucidate the challenges faced by autoregressive models in specific scenarios. Through this analysis, we introduce the concept of subgoal imbalance—wherein some subgoals are inherently more difficult than others—which offers insights into these difficulties. We then extend our discussion in §3.2 to examine how diffusion models can more effectively address and learn these hard subgoals, effectively overcoming the limitations observed in autoregressive approaches. We finally propose Multi-Granularity Diffusion Modeling (MGDM; §3.3) as a natural extension of discrete diffusion models to better address these challenges and improve performance on complex tasks requiring planning and reasoning.
@@ 3.1 Subgoal Imbalance in Autoregressive and Diffusion Modeling
(Subgoal imbalance due to the unknown data distribution $q({\mathbf{x}})$) Given the data ${\mathbf{x}}$ sampled from an unknown data distribution $q({\mathbf{x}})$, the difficulty of learning each subgoal ${\bm{x}}_{n}$ can differ significantly based on how we parametrize the model distribution, and some subgoals may require substantially more data to learn or may even be infeasible to learn.
Subgoal imbalance in autoregressive modeling.
Given the data ${\mathbf{x}}$ sampled from an unknown data distribution $q({\mathbf{x}})$, the naive autoregressive modeling parametrizes the model distribution $p_{\bm{\theta}}({\mathbf{x}})$ as $p_{\bm{\theta}}({\bm{x}}_{1})\prod_{n=2}^{N}p_{\bm{\theta}}({\bm{x}}_{n}\mid{\bm{x}}_{1:n-1})$, the difficulty of learning each subgoal ${\bm{x}}_{n}$ can differ significantly as given only the left context, and some subgoals may require substantially more data to learn or may even be infeasible to learn.
@@ 3.2 Effective Hard Subgoal Learning in Diffusion Modeling
Further analysis of the unweighted loss $u({\mathbf{x}}_{0},{\mathbf{x}}_{t},n;{\bm{\theta}})$ in the diffusion model, based on 1,000 samples of ${\mathbf{x}}_{t}\sim q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$, reveals a clear trend: as the number of timesteps increases, resulting in more noise in ${\mathbf{x}}_{t}$, objectives in smaller timesteps (i.e., recovery from less noisy data) become significantly easier to learn. From a multi-view learning perspective (Xu et al., 2013), each ${\mathbf{x}}_{t}$ can be interpreted as a distinct view of ${\mathbf{x}}_{0}$, where each view provides different information about ${\mathbf{x}}_{0}$. In the diffusion process, by exploring the consistency and complementary properties of different views offered by a diverse range of interrelated objectives $u({\mathbf{x}}_{0},{\mathbf{x}}_{t},n;{\bm{\theta}})$, our findings suggest that objectives challenging to learn in AR models become more effective, promising, and exhibit better generalization in diffusion models.
@@ 3.3 Multi-granularity Diffusion Modeling
These observations provide valuable insights, i.e., diffusion modeling builds on a diverse range of interrelated views from the data ${\mathbf{x}}_{0}$ to handle a challenging subgoal. To handle multiple challenging subgoals in real data, we should prioritize different subgoals based on their difficulty during the learning process to achieve more effective learning outcomes and faster convergence, and this naturally translates to prioritizing difficult views as the learning of a subgoal depends on learning interrelated views related to it. Building on this, we propose the multi-granularity diffusion model as a natural extension of the discrete diffusion model.
@@
For a sequence of length $N$, the probability of sampling the same ${\mathbf{x}}_{t}$ in AR is 1. However, in diffusion, this probability reduces to $1/C_{N-1}^{t(N-1)/T}$ due to the randomness in sampling ${\mathbf{x}}_{t}$, potentially reducing the training efficiency of diffusion models. We note that Equation (7) employs a sequence-level reweighting term $w(t)$ to indicate the importance of ${\mathbf{x}}_{t}$. However, individual tokens within the sequence, given their imbalanced difficulties, are not properly reweighted. To address this, we propose multi-granularity diffusion modeling (MGDM), which introduces an additional token-level reweighting mechanism to enhance training efficiency:
$L_{\mathrm{\textsc{MGDM}}}=\sum_{n=1}^{N}\sum_{t=1}^{T}w(t)v({\mathbf{x}}_{t,n})u({\mathbf{x}}_{0},{\mathbf{x}}_{t},n;{\bm{\theta}}),$ (8)
@@ 4.1 Countdown
MGDM 76.0 x1
As shown in Table 4.1, the performance of the GPT-4 with IO, CoT, and CoT-SC prompting methods from  Yao et al. (2024) is unsatisfactory for the given task, with only accuracy below 10%. The introduction of ToT, which incorporates a search algorithm designed by human experts into the decoding process, significantly enhances the performance of GPT-4. This integration allows the AR model to backtrack as needed, resulting in notable improvements. However, this paradigm requires the assessment of intermediate steps using LLM, resulting in considerable computational costs due to the need for multiple LLM calls. We list the token cost in Table 4.1 with more details in Appendix D.1. ToT consumes 186 times more tokens than MGDM, showcasing the ‘internal’ search capability by promoting global consistency in diffusion modeling. In summary, our model, despite having a parameter size of only 85M, significantly outperforms both the AR task-specific model of the same size (GPT-2 Scratch) in terms of performance and the larger general pre-trained model (GPT-4) in computation cost, indicating it is challenging for model scaling and decoding strategies to substitute the advantages of modeling paradigm.
@@ 4.3 Boolean Satisfiability Problem
As shown in the right figure of Figure 4, MGDM performs well in solving scenarios with five variables, while the AR model falls slightly short. As the number of variables increases, both our model and the AR model experience a certain degree of decrease in accuracy. However, the performance gap between the two models widens as the difficulty of the task increases. This indicates that our diffusion model exhibits a more pronounced advantage in handling more challenging tasks than the AR counterpart.
@@ 4.4 Analysis
Figure 5: Evaluation accuracy throughout the training process for AR and MGDM with different reweighting strategies.
@@
We assess the trade-off between accuracy and decoding speed by comparing the performance of the AR model (GPT-2 Scratch 85M) and MGDM (85M). The speed metric is determined by the number of samples processed using a batch size of 1 on the NVIDIA GeForce V100 GPU. As shown in Figure 6(a), MGDM can flexibly control the trade-off between accuracy and decoding speed by varying the diffusion timesteps. Notably, by employing just one diffusion step, MGDM demonstrates a remarkable 10x improvement in speed compared to AR, while maintaining superior accuracy with 75% and 12.7% compared to 45.8% and 5.1% of AR on Countdown 4 and 5, respectively. We observed that the slope of countdown 4 is smaller compared to countdown 5 in the trade-off. This suggests that for tasks with lower complexity, diffusion demonstrates a more noticeable speed advantage by setting a smaller diffusion step. In addition, it also indicates that sacrificing some efficiency for performance improvements becomes particularly evident when dealing with more intricate tasks.
@@
To gain a deeper understanding of error patterns in AR and MGDM, we conducted an error analysis on Countdown 4. For instance, given the input “7,38,3,1” and the target number 14, a correct solution would be “97-38=59,59-17=42,42/3=14”. We divide the solution into four parts: equation 1, equation 2, equation 3, and answer checking. First, from a calculation perspective, we assess the error ratio in each equation by comparing the left-hand side and the right-hand side, regardless of whether the correct number was chosen. As shown in Figure 6(b), the bar plot demonstrates that the majority of calculation errors for AR are concentrated in Equation 3. For example, given input “16,4,40,51” and target 87, the prediction of AR is “51-40=11,16*4=64,11+64=87” while the correct solution is “16/4=4,40+51=91,91-4=87”. The first two equations are calculated correctly, but the last one is incorrectly forced to equal 87. Moreover, in many cases, the error rate for the last equation in AR (48.9%) is significantly higher than that of the previous equations (0.2% for the first and 7.2% for the second). This indicates that the model’s lack of planning capability results in its realization of being unable to achieve the goal only at the end. By that point, it cannot correct the previous errors, leading it “regretfully” to rely on incorrect equations to reach the goal (e.g., the number 87 in the above case). This significantly increases the number of calculation errors in the third equation. We call this phenomenon ‘The Regretful Compromise’. The reason for this is that the AR model made incorrect choices of numbers or operations in previous equations. This is demonstrated by examining the step at which the models fail the task, as depicted in Figure 6(b). It is evident that there is a notable frequency of planning errors in the first equation for the AR model, where the number of errors is significantly higher compared to our model. This highlights the limitations of the left-to-right decoding approach in AR, which adversely affects its planning ability.
@@ 5.2 Non-autoregressive Modeling
The non-autoregressive (NAR) generation method, which produces all target tokens simultaneously given the source context, is first proposed by  (Gu et al., 2017) in the text field for machine translation, primarily due to the efficiency consideration. While a series of advancements have been made afterward (Lee et al., 2018; Gu et al., 2019; Ghazvininejad et al., 2019; Qian et al., 2021; Huang et al., 2022, inter alia), traditional NAR models still tend to underperform AR models in terms of generation quality (Xiao et al., 2023). Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), a powerful class of generative models known for their impressive image-generation capabilities (Dhariwal & Nichol, 2021), have recently been applied to the field of text generation (Hoogeboom et al., 2021; Austin et al., 2021; Li et al., 2022; Campbell et al., 2022; Dieleman et al., 2022; Chen et al., 2023; Ye et al., 2023b; Lovelace et al., 2024), reinforcement learning (Janner et al., 2022; Chi et al., 2023) and protein design (Xu et al., 2022; Hoogeboom et al., 2022b; Corso et al., 2023). In essence, diffusion models perform a multi-step denoising process to progressively convert a random noise into a data sample, and the denoising procedure can be seen as parameterizing the gradients of the data distribution (Song & Ermon, 2019), connecting them to score matching (Hyvärinen & Dayan, 2005) and energy-based models (LeCun et al., 2006). For text, the diffusion model can be seen as an extension of the traditional iterative NAR models (Gong et al., 2022) and has been shown to approach or outperform AR models on text perplexity (Han et al., 2023; Lou et al., 2023; Gulrajani & Hashimoto, 2024), diversity (Gong et al., 2022; 2023; Zhang et al., 2023) as well as various seq-to-seq tasks (Wu et al., 2023b; Zheng et al., 2023; Ye et al., 2024). In this work, we compare diffusion with AR from a perspective of subgoal imbalance and demonstrates the effectiveness of diffusion in tasks requiring complex reasoning and planning.
@@ 6 Conclusion
This paper presents an extensive analysis of the limitations of auto-regressive (AR) language models when applied to planning tasks that involve deliberate planning, both in controlled settings and real-world contexts. Based on an advanced understanding, we propose an improved diffusion model, named MGDM, that performs significantly better than AR and previous diffusion models on various sophisticated planning tasks. Our findings underscore the necessity to reevaluate the sequence modeling paradigm for modern large language models, especially in tackling challenging problem-solving tasks.
Acknowledgments
This research was supported in part by the joint research scheme of the National Natural Science Foundation of China (NSFC) and the Research Grants Council (RGC) under grant number N_HKU714/21.